OpenStudy (anonymous):

Identify the correct x intercept and y intercept of the equation x + 7y = -21. x intercept = (3, 0); y intercept = (0, 21) x intercept = (-3, 0); y intercept = (0, -21) x intercept = (21, 0); y intercept = (0, 3) x intercept = (-21, 0); y intercept = (0, -3)

OpenStudy (anonymous):

x intercept is -21 and y intercept is-3

OpenStudy (anonymous):

to find x intercept put y =0 and to find y interceput put x=0 ... simply

OpenStudy (anonymous):

What about Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. y + 3 = -6(x + 2) y + 2 = 6(x + 3) y - 3 = -6(x - 2) y - 2 = 6(x - 3)

OpenStudy (anonymous):

1st option

OpenStudy (anonymous):

so simple take (y-y1)=m(x-x1) (x1,y1) is given point so 1st

OpenStudy (anonymous):

Thank You so much!

OpenStudy (anonymous):

Which of the following is an example of positive correlation? House address vs math test score Age vs commute time to school Time worked vs money earned Time eating vs food left

terenzreignz (terenzreignz):

Anytime you need to find the equation of a line with slope \(\color{green}m\) and passing through the point \((\color{red}a \ , \color{blue}b)\) You use this form: \[\Large y - \color{blue}b = \color{green}m(x-\color{red}a)\] This is basically what @khurramshahzad said already... just clarifying... And you might want to give a (more thorough) explanation first before jumping into the answer... just saying...

OpenStudy (anonymous):

ok bro

terenzreignz (terenzreignz):

As for that question, two values would be in positive correlation IF when one increases, the other would also increase, and when one decreases, the other also decreases. Which of those pairs most satisfactorily fulfills this criterion?

OpenStudy (anonymous):

Time worked vs money earned?

terenzreignz (terenzreignz):

The more time you work, the more money you earn? You wish ;) JK yeah, that's it :)

OpenStudy (anonymous):

Haha Thanks! (: And good one! Lol

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

How would I figure out this? Wake Up coffee shop offers two different sizes of coffee. The small coffee costs $2 and the grande coffee costs $3. It is your turn to buy coffee for the office and the bill came to $18. This scenario can be modeled by the equation 2s + 3g = 18. Which is not a possible combination of small and grande coffees?

terenzreignz (terenzreignz):

Okay... you haven't posted choices... I'll assume they're of the form... s small coffee's g grandes or words to that effect. What you do is replace the s and g (respectively) in \[\large 2\color{red}s + 3\color{blue}g = 18\] with the value in each choice and see if it adds up. (ie, the left-side would add up to 18) So, say a choice goes 3 small's 4 grande's Well.. \[\Large 2(\color{red}3) + 3(\color{blue}4)= 6 + 12 = 18 \] So it adds up, as it turned out to be equal to 18. So s = 3 and g = 4 is in fact a possible combination. Your task is to find the choice in which when you replace s and g accordingly, 2s + 3g does not add up to 18.

OpenStudy (anonymous):

6 small and 2 grande

terenzreignz (terenzreignz):

Let's check :) What's \[\Large 2(\color{red}6) + 3(\color{blue}2) = ?\]

OpenStudy (anonymous):

18 (:

terenzreignz (terenzreignz):

Well, don't be so smiley about it >:) You're supposed to be finding the pair that DON'T add up to 18 when you sub...

OpenStudy (anonymous):

its 18 ... if u take s=2 then the conbination will never give u a natural number for g means thats the answer taking s=2

terenzreignz (terenzreignz):

I quote your problem... "Which is not a possible combination of small and grande coffees? "

OpenStudy (anonymous):

9 small and 1 grande

terenzreignz (terenzreignz):

Check? \[\Large 2(\color{red}9) + 3(\color{blue}1) = ?\]

OpenStudy (anonymous):

can u revise ur question to clear ur mind ... i said any value that does`nt satisfy ur given equation will b ur answer as ur theme or question states

OpenStudy (anonymous):

So that would come out to 21

terenzreignz (terenzreignz):

there are choices @khurramshahzad They just weren't posted :) And that's correct, @Hopers15 so are you now confident with your answer? :)

OpenStudy (anonymous):

Yes, Thank You (:

OpenStudy (anonymous):

On this one first how would I indentfy the slope? ]Identify the correct slope and y intercept of the equation 6x - 3y = 12. slope = 2; y intercept at (0, -4) slope = -2; y intercept at (0, 4) slope = one half; y intercept at (0, 15) slope = one half; y intercept at (0, 9)

terenzreignz (terenzreignz):

Okay, we better clear up a little slope-catching tricks, for your convenience... I hope you don't mind if I put up a general case... \[\Large \color{red}px + \color{blue}qy =c\] And how to find the slope? first, can you put the equation in this form... \[\Large y = \color{green}mx + \color{orange}b\]?

OpenStudy (anonymous):

Okay

terenzreignz (terenzreignz):

I have to go now... I'm so sorry, @khurramshahzad could you please take over? I'm feeling very sleepy now~ won't be able to help as effectively... ----------------------------------- Terence out

OpenStudy (anonymous):

ok